OpenStudy (anonymous):

Find f(a), f(a + h), and the difference quotient (f(a + h) − f(a))/h where h ≠ 0. f(x) = 7 − 4x + 3x2

OpenStudy (anonymous):

\[f(a+h) = 7-4(a+h)+3(a+h)^{2}\] \[f(a+h) - f(a) = 7-4(a+h)+3(a+h)^{2} - (7-4a+3a ^{2})\] = \[6ah + 3h ^{2} - 4h\] After dividing out by h, you should get 6a - 4 + 3h